Lyn gave half of his salary and 500 to his parents .then he spent half of his remaining salary and 1000 on rent. How much was his salary if he had 5600 left?
Found 2 solutions by mathdad1965, ikleyn:
Answer by mathdad1965(8)

(1/2)x + 500 + (1/2)x + 1000 = 5600
Solve for x to find your answer.

Answer by ikleyn(52775)

.
Lyn gave half of his salary and 500 to his parents . then he spent half of his remaining salary and 1000 on rent.
How much was his salary if he had 5600 left?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

0.5x + 500 to his parent.


Remaining is x - (0.5x + 500) = 0.5x - 500


The equation is 


x - (0.5x + 500) - 0.5*(0.5x - 500) - 1000 = 5600


Simplify and solve for x.

The solution by the tutor "mathdad1965" is wrong.